OpenStudy (anonymous):

how can i calculate the charge on a point q, given the net electric force and the values of 2 other charges?

OpenStudy (anonymous):

you've got to show the whole question mam

OpenStudy (nottim):

You should add the given info into a diagram and/or redraw the diagram.

OpenStudy (anonymous):

maybe this will help ;)

OpenStudy (nottim):

Yup. I can't really much though. I can't remember much from this,=.

OpenStudy (nottim):

Eashmore, can you do anything here?

OpenStudy (anonymous):

I can do it.

OpenStudy (anonymous):

We have to balance forces. We know the expression for electromagnetic force is\[F_E = k_e {q_1 q_2 \over r^2}\] First, let's find the component of the net force that acts in the x and y-directions. \[F_{x,3} = F_3 \cos(\theta)\]\[F_{y,3} = F_3 \sin(\theta)\] We note that both forces act positive in their respective directions. Since \(q_1\) is positive and located to the left of \(q_3\), \(q_3\) must be positive if the force from \(q_1\) on \(q_3\) is to be in the positive x-direction. The same thought process verifies this when relating \(q_2\) and \(q_3\). Additionally, let's note that the force from \(q_1\) acts on \(q_3\) solely in the x-direction, and \(q_2\) acts on \(q_3\) solely in the y-direction. Let's balance the forces in x and y-directions now. \[F_{x,3} = k_e {q_1 q_3 \over r^2}\]and\[F_{y,3} = k_e {q_2 q_3 \over r^2}\]

OpenStudy (anonymous):

ok this makes sense so far

OpenStudy (anonymous):

You should be able to rearrange the last two equations for \(q_3\). You should get the same value.

OpenStudy (anonymous):

what do you mean? set them equal? or factor q3 out of the equation?

OpenStudy (anonymous):

Pick one and solve it for \(q_3\). If we solved both for \(q_3\), both equations should produce the same value for \(q_3\).

OpenStudy (anonymous):

Fx^2 + Fy^2 = 30.187 ^2 ? and sub both equations in for Fx and Fy

OpenStudy (anonymous):

I have two unknowns in the above equations yes? so how do I solve for Q3?

OpenStudy (anonymous):

factoring?

OpenStudy (jamesj):

You only have one unknown: \( q_3 \). You know q1 and q2; you know r for each equation; and you know the constant \( k_e \). So you only need to solve for q3.

OpenStudy (jamesj):

(Notice the r=1 m for one of the forces, and r=2m for another force)

OpenStudy (jamesj):

In other words, for both of the equations .... the only variable you don't know is q3. So you can use either equation to solve for q3.

OpenStudy (jamesj):

talk to me ... what's not making sense?

OpenStudy (jamesj):

ok ...I'm out here of here then.

OpenStudy (anonymous):

i do not know Fx or Fy I only know the sum of both..sorry having internet problems

OpenStudy (anonymous):

You know \(F_x\) and \(F_y\). Take the total force as being \(F_3\), then use the trigonometric relations I gave you for \(F_x\) and \(F_y\). \[F_x = F_3 \cos(\theta) ~ {\rm and} ~ F_y = F_3 \sin(\theta)\]